Emma has had a puncture and the air is leaking from the tire. In addition, it is starting to rain. From the beginning, the tire held 5 dl of air with a pressure of 300 kPa. We can assume that the temperature is constant and that the average pressure during the leak is 165 kPa

Is any heat added to the leaking air? I want a motivation to the answer.
1
Expert's answer
2021-03-14T19:20:42-0400
"(P_i-P_a)V_i=(P_f-P_a)V_f\\(300-101.325)5=(165-101.325)V_f\\V_f=15.6\ dL=1.56\ L"